Calculate the size of an organism that occupies 26 um spaces under scanner, LPO and HPO using stage micrometer and ocular micrometer if the calibration factor is 0.65, 0.75 and 0.85
1
Expert's answer
2020-09-18T12:48:32-0400

Total magnification"=4\times10=40×"

Occupied space"=26\mu m"

Calibration factor"=0.65"

size of organism"=\frac{26\mu m}{40\times 0.65}=1\mu m"


calibration factor="0.75"

size of organism"=\frac{26\mu m}{40\times0.75}=0.87\mu m"


Calibration factor"=0.85"

size of organism"=\frac{0.85\mu m}{40\times0.85}=0.765"


For low power(L.P.O)="Magnification=10\times10=100\times"

calibration factor="0.65;" size of organism ="\frac{26\mu m}{100\times0.65}=0.4\mu m"


calibration factor="0.75" ;size of organism="\frac{26\mu m}{100\times0.75}=0.35\mu m"


calibration factor="0.85" ; "\frac{26\mu m}{100\times0.85}=0.31\mu m"


For high power(H.P.O)=magnification="40\times10=400\times"

calibration="0.65;" size of organism="\frac{26\mu m}{400\times0.65}=0.1\mu m"


Calibration="0.75;" size of organism="\frac{26\mu m}{400\times0.75}=0.087\mu m"


calibration="0.85;" size of organism="\frac{26\mu m}{400\times0.85}=0.077\mu m"